OpenStudy (anonymous):

parents need $95000 in 10 years for college expenses, and that the bank account earns 11.75% compounded continuously. Round your answers to the nearest cent. (You may need to compute your answers to 4 or more decimal places before you round to the nearest cent.) ) At what constant, continuous rate must the parents deposit money into the account in order to save the money? b) If the parents instead deposit a lump sum now, how much must the deposit be to attain the goal?

OpenStudy (anonymous):

A=P(1+R%)^10 95000=P(1+11.75%)^10 95000=P(111.75/100)^10 95000=P(3.0372) 95000/3.0372=P 31,278.67066=P FOR 10 YEARS MAYBE P/10

OpenStudy (anonymous):

would you like to help with two more?

OpenStudy (anonymous):

was this right ???

OpenStudy (anonymous):

nope haha

OpenStudy (anonymous):

Contiuous is Pe^rt

OpenStudy (anonymous):

D- Deposit r- rate Let assume that we are depositing once a year for now 1st Year = (D*E^(RT)) 2nd Year = (D+( D*E^RT))*E^(RT)=DE^rt+DE^(rt)^2 3rd Year = DE^rt+DE^(rt)^2+DE^(rt)^3 10 Year = DE^rt+DE^(rt)^2+DE^(rt)^3+......+DE^(rt)^10